OpenStudy (kydawg):

Rewrite with only sin x and cos x. sin 3x

jhonyy9 (jhonyy9):

sin3x = sin(2x+x) = ? use formzula sin(a+b) = ?

OpenStudy (mathstudent55):

\(\sin (x+y) = \sin x \cos y + \cos x \sin y\) \(\sin 3x=\) \(= \sin (x + 2x)\) \(=\sin (x+2x) = \sin x \cos 2x + \cos x \sin 2x\) Now use the sine of 2x and the cosine of 2x to simplify further.

OpenStudy (kydawg):

The answer is 2cos^2x sinx + sinx - 2sin^3x?

OpenStudy (mathstudent55):

Also, \(\cos (2x) = \cos^2 x - \sin ^2 x\) \(=\sin x \cos 2x + \cos x \sin 2x\) \(=\sin x (\cos^2 x - \sin ^2 x) + 2\cos x \sin x \cos x\) \(=\sin x \cos^2 x - \sin^3 x + 2 \sin x \cos^2 x \) \(= 3\sin x \cos^2 x - \sin^3 x\)

OpenStudy (mathstudent55):

You are also correct.

OpenStudy (kydawg):

Thanks.

OpenStudy (mathstudent55):

yw